OpenStudy (anonymous):

*MEDAL* Solve And Explain Solve by substitution. Show each step of your work. 2x + y = 7 3x + 5y = 14

OpenStudy (anonymous):

2x+y=7 3x+5y = 14 y= 7-2x 3x + 5(7 - 2x) = 14 3x + 35 - 10x = 14 solve for x

OpenStudy (anonymous):

Oh dang its a competition between Snagya and CWRW!!!

OpenStudy (anonymous):

cwrw is that answer true check it

OpenStudy (anonymous):

and btw if its true I have 2 more questions lol

OpenStudy (anonymous):

Solve by multiplication and addition. Show each step of your work. 4x – 3y = 11 3x + y = 18

OpenStudy (anonymous):

was that for the first or second?

OpenStudy (cwrw238):

-what sangya21 has posted is correct

OpenStudy (cwrw238):

I'm having trouble with my computer

OpenStudy (anonymous):

(4x - 3y = 11)(1) (3x + y = 18) (3) (4x -3y) + (6x + 3y) 11 + 54 solve for x

OpenStudy (anonymous):

last question sangya and you get medal

OpenStudy (anonymous):

okk

OpenStudy (anonymous):

Solve the following word problem using two variables. Show each step of your work. Tanya is five years younger than Heidi. The sum of their ages is 23. Find both ages.

OpenStudy (anonymous):

let age of Tanya be x and age of heidi be y y = x+5 x+ y = 23 do the substitution method

OpenStudy (anonymous):

can you do the substitution method for meh pwease

OpenStudy (anonymous):

x + x + 5 = 23

OpenStudy (anonymous):

thx and the winner is...

OpenStudy (anonymous):

cwrw

OpenStudy (anonymous):

jk jk

OpenStudy (cwrw238):

lol

OpenStudy (anonymous):

cwrw you were having problems with your computer so I am sorry

OpenStudy (anonymous):

thanks sangya

OpenStudy (anonymous):

Anytime.

OpenStudy (anonymous):

I also fanned you so we can stay in contact (: thank you again

OpenStudy (anonymous):

Sure. :)